# Private Frequency Estimation Via Residue Number Systems

### v1

@@ 1 Introduction
Together, these four dimensions form a multi-bottleneck regime: in large-scale telemetry and federated analytics, client bandwidth, server compute, statistical accuracy, and privacy risk may each become the dominant constraint depending on the deployment.
@@
This work addresses this multi-bottleneck challenge when designing efficient mechanisms for federated-analytics deployments that require locally differentially private frequency estimation over a finite domain $k]=\{0,\dots,k-1\}$. In this setting, each user holds a private input $x_{i}$ and the goal is for an untrusted server to recover an accurate estimate of the population histogram $\mathbf{f}\in\mathbb{R}^{k}$, where $f_{v}=\tfrac{\#\{i:x_{i}=v\}}{n}$. After collecting $n$ randomized reports $\{Y_{i}\}_{i=1}^{n}$, the server computes an estimate $\hat{\mathbf{f}}$ aiming to minimize its distance from $\mathbf{f}$ under some norm $\lVert\mathbf{f}-\hat{\mathbf{f}}\rVert$. In line with prior literature (feldman2022; kairouz2016discrete; tianhao2017; Hadamard), we quantify estimation error using the expected $\ell_{2}$ norm, and focus on the mean squared error (MSE) metric: $\mathrm{MSE}\;=\;\frac{1}{k}\,\operatorname{E}\bigl[\lVert\hat{\mathbf{f}}-\mathbf{f}\rVert_{2}^{2}\bigr]$.
@@ 4 Analysis of MSS
This section analyzes MSS along the four axes of the multi-bottleneck regime highlighted in Section 1. We bound its communication and decoding costs, derive closed-form and worst-case MSE expressions in terms of the condition number $\kappa$, show how moduli selection controls $\kappa$, and quantify its resilience to data reconstruction attacks.
@@ 5 Experimental Results
We benchmark our ModularSubsetSelection against state-of-the-art single-message frequency oracles: ProjectiveGeometryResponse, RandomizedResponse, SubsetSelection, and OptimalUnaryEncoding (the optimized RAPPOR variant). Since worst-case MSE is distribution-independent (Table 1), we adopt the synthetic Zipf ($s=3$) and Spike ($\mathbf{f}=[1,0,\ldots,0]$) benchmarks from (feldman2022), both known to induce high estimation variance. Unless noted otherwise, we fix n=10,000n=10{,}000 users, domain $k\in\{1024,22000\}$, privacy budget $\varepsilon\in\{0.5,1.0,\dots,4.5,5.0\}$, and average results over 300 independent trials.
@@
Fig. 2 shows the number of bits each user must transmit under both SS and MSS, as a function of $\varepsilon$ for $k=1024$ and k=22,000k=22{,}000. Since MSS relies on a randomized moduli selection process, we report its average and standard deviation over the 300 runs. Across all settings, MSS consistently achieves lower communication cost than SS, up to one-half in high-privacy regimes, while retaining comparable accuracy (see Fig. 1). We omit GRR and PGR from the figure for clarity: their per-report message length is fixed for a given $k$ (both use $O(\log k)$ bits) and is already summarized analytically in Table 1. GRR and PGR therefore form communication-efficient baselines, but as shown in Fig. 1 and Table 2, they pay respectively in much higher MSE (GRR) or substantially higher decoding cost (PGR).
(a) $k=1024$
@@
We now compare the server-side runtime of our MSS protocol against the state-of-the-art PGR scheme by feldman2022. We run both protocols on the Zipf dataset of size n=10,000n=10{,}000 and domain size k=22,000k=22{,}000, across several privacy levels. Table 2 reports the average and standard deviation of the server decoding time (in seconds) over 300 trials. MSS consistently outperforms PGR by large margins, achieving decoding speed-ups between $11\times$ and $448\times$. This performance gap stems from their algorithmic differences: MSS solves a sparse weighted least-squares problem, while PGR relies on algebraic decoding over finite fields. We do not plot GRR here, since its server cost is essentially a single histogram pass $O(n+k)$ and thus serves as a trivial lower bound on runtime; however, as Fig. 1 and Fig. 3 show, GRR is not competitive in our multi-bottleneck regime due to its much worse utility and attackability. The runtime spike at $\varepsilon=4.5$ for PGR likely arises from parameter rounding and structural constraints in its projective geometry design.
@@
We now evaluate the vulnerability of each LDP protocol to a single-message data reconstruction attack (DRA) (see Section 4.4). Fig. 3 shows the empirical DRA as a function of the privacy budget $\varepsilon$, under the Zipf distribution for domain sizes $k=100$ and $k=1024$. MSS consistently achieves the lowest DRA across all $\varepsilon$ values, confirming its robustness to reconstruction attacks. This is due to its modular randomization strategy, which distributes the probability mass across multiple residue classes, making inference more challenging. In contrast, GRR and SS exhibit higher attackability, especially for small $k$, as their output space is tightly linked to the input domain. PGR behaves comparably to SS/MSS/OUE at moderate $\varepsilon$, but for larger budgets its DRA increases sharply when $k$ is smaller than the projective-domain size $K=(q^{t}-1)/(q-1)$ required by its internal geometry. This truncation mismatch breaks PGR’s combinatorial symmetry and makes certain messages disproportionately informative. A fair comparison using the non-truncated setting $k=K$ is provided in Appendix D.
@@ D Complementary Results
Fig. 4 compares the analytical MSE derived for both SS and MSS against their empirical counterparts, under both Zipf and Spike distributions for $k\in{1024,22{,}000}$. The analytical expressions closely match the empirical observations across all regimes of $\varepsilon$, consistently tracking the empirical trends. This validates the tightness and reliability of our closed-form MSE derivation for MSS (Eq. (3)) in practice, regardless of the underlying data distribution or domain size. The results provide strong evidence that our theoretical analysis generalizes well and can be used for practical design decisions without the need for repeated empirical tuning.
@@
Figure 4: Comparison between analytical and empirical MSE for SS and MSS protocols, across a range of privacy budgets $\varepsilon$, for $k=1024$ and k=22,000k=22{,}000 under both Zipf and Spike distributions. Each empirical MSE is averaged over 300 runs, while analytical curves are computed in closed-form expressions.
@@
Fig. 5 presents an ablation study analyzing how the performance of the MSS protocol varies with fixed numbers of moduli $\ell\in\{3,6,9,12,15\}$, compared to the analytically optimized MSS[OPT]. We observe that the relationship between $\ell$, utility (MSE), and communication cost (bit cost) is non-linear: in some regimes, intermediate values such as $\ell=9$ can outperform both smaller ($\ell=3$) and larger ($\ell=12$) settings in terms of accuracy and communication efficiency. This highlights the complex trade-offs induced by modular encoding, where increasing $\ell$ does not guarantee monotonic improvements. Notably, MSS[OPT] consistently selects a configuration that achieves near-optimal utility across all privacy budgets, confirming the effectiveness of our moduli selection strategy. It is important to emphasize that our optimization procedure (Section 4.3 and Section A) currently targets minimizing analytical MSE only. This objective implicitly balances communication and robustness in many settings, but it is not guaranteed to yield optimal trade-offs across all criteria. Future work could extend this framework to support multi-objective optimization, for instance, incorporating additional metrics such as bit cost or attackability, enabling more fine-grained control over privacy-utility-efficiency trade-offs.
Figure 5: Ablation study showing the impact of the number of moduli $\ell\in\{3,6,9,12,15\}$ on the utility (left), communication cost (middle), and attackability (right) of the MSS protocol. The dashed black curve represents the performance of our ModularSubsetSelection protocl (i.e., MSS[OPT]), which automatically selects $\ell$ and the moduli via our analytical optimization procedure. Results are averaged over 300 runs for $k=1024$, under the Zipf distribution.
@@
To evaluate the tightness of our analytical DRA derivations (Section 4.4 and Section B), we compare theoretical and empirical data reconstruction attack (DRA) for both SS and MSS across privacy budgets. As shown in Fig.6, SS exhibits a near-perfect match between analytical and empirical ASR. For MSS, however, the analytical ASR bound consistently overestimates the true attackability. This gap is theoretically expected: the analytical expression used in the main paper is a conservative upper bound derived using Jensen’s inequality over the random support size of the modular posterior (see Section B). The underlying randomness of the filler set in each residue block, and the multiplicity of values per residue, makes the exact computation of MSS’s ASR more intricate, resulting in a looser but guaranteed-safe upper bound. These results show that while the MSS ASR bound is not tight, it still provides a safe analytical proxy and reinforces that MSS offers strong practical protection against reconstruction attacks.
Figure 6: Empirical vs. analytical data reconstruction attack (DRA) under Zipf and Spike distributions, for both small and large domains. While $S$ closely matches its analytical ASR, MSS shows a consistent gap, confirming that the bound is conservative.
@@
While these trends hold generally, it is important to highlight a distinctive behavior of PGR: unlike other mechanisms, PGR’s internal domain size is determined by the projective geometry induced by the privacy budget $K(\varepsilon)=(q^{t}-1)/(q-1)$ with $q\approx e^{\varepsilon}+1$. To isolate the impact of this dependence, we perform an additional ablation (see Fig. 9) in which, for each privacy level $\varepsilon$, we set the evaluation domain to the corresponding projective size $K(\varepsilon)$ (computed using a reference scale $\mathrm{BASE\_K}=100$), and we run all protocols on this same domain. This removes the truncation mismatch that affects PGR when $k<K(\varepsilon)$ and places all baselines in the geometry naturally required by PGR. As shown in Fig. 9, once this geometry-aligned setting is used, PGR’s attackability no longer exhibits the sharp rise observed in Figs. 3 and 8, while MSS continues to provide the lowest DRA across all $\varepsilon$.
@@
Figure 9: Geometry-aligned ablation for PGR. For each privacy level $\varepsilon$, the domain size is set to the projective dimension $K(\varepsilon)$ naturally induced by PGR (using $\mathrm{base_{k}}=100$), and all protocols are evaluated on this shared domain. This removes the truncation mismatch present in prior experiments, eliminating PGR’s high-$\varepsilon$ spike. MSS remains the most robust to reconstruction attacks.

### v2

@@ 1 Introduction
Together, these four dimensions form a multi-constraint regime: in large-scale telemetry and federated analytics, client bandwidth, server compute, statistical accuracy, and privacy risk may each become the dominant constraint depending on the deployment.
@@
This work addresses this multi-constraint challenge when designing efficient mechanisms for federated-analytics deployments that require locally differentially private frequency estimation over a finite domain $k]=\{0,\dots,k-1\}$. In this setting, each user holds a private input $x_{i}$ and the goal is for an untrusted server to recover an accurate estimate of the population histogram $\mathbf{f}\in\mathbb{R}^{k}$, where $f_{v}=\tfrac{\#\{i:x_{i}=v\}}{n}$. After collecting $n$ randomized reports $\{Y_{i}\}_{i=1}^{n}$, the server computes an estimate $\hat{\mathbf{f}}$ aiming to minimize its distance from $\mathbf{f}$ under some norm $\lVert\mathbf{f}-\hat{\mathbf{f}}\rVert$. In line with prior literature (feldman2022; kairouz2016discrete; tianhao2017; Hadamard), we quantify estimation error using the expected $\ell_{2}$ norm, and focus on the mean squared error (MSE) metric: $\mathrm{MSE}\;=\;\frac{1}{k}\,\operatorname{E}\bigl[\lVert\hat{\mathbf{f}}-\mathbf{f}\rVert_{2}^{2}\bigr]$.
@@ 4 Analysis of MSS
This section analyzes MSS along the four axes of the multi-constraint regime highlighted in Section 1. We bound its communication and decoding costs, derive closed-form and worst-case MSE expressions in terms of the condition number $\kappa$, show how moduli selection controls $\kappa$, and quantify its resilience to data reconstruction attacks.
@@ 5 Experimental Results
We benchmark our ModularSubsetSelection against state-of-the-art single-message frequency oracles: ProjectiveGeometryResponse, RandomizedResponse, SubsetSelection, and OptimalUnaryEncoding (the optimized RAPPOR variant). Since worst-case MSE is distribution-independent (Table 1), we adopt the synthetic Zipf ($s=3$) and Spike ($\mathbf{f}=[1,0,\ldots,0]$) benchmarks from (feldman2022), both known to induce high estimation variance. Unless noted otherwise, we fix n=10,000n=10{,}000 users, domain $k\in\{1{,}024,22{,}000\}$, privacy budget $\varepsilon\in\{0.5,1.0,\dots,4.5,5.0\}$, and average results over 300 independent trials.
@@
Fig. 2 shows the number of bits each user must transmit under both SS and MSS, as a function of $\varepsilon$ for k=1,024k=1{,}024 and k=22,000k=22{,}000. Since MSS relies on a randomized moduli selection process, we report its average and standard deviation over the 300 runs. Across all settings, MSS consistently achieves lower communication cost than SS, up to one-half in high-privacy regimes, while retaining comparable accuracy (see Fig. 1). We omit GRR and PGR from the figure for clarity: their per-report message length is fixed for a given $k$ (both use $O(\log k)$ bits) and is already summarized analytically in Table 1. GRR and PGR therefore form communication-efficient baselines, but as shown in Fig. 1 and Table 2, they pay respectively in much higher MSE (GRR) or substantially higher decoding cost (PGR).
(a) k=1,024k=1{,}024
@@
We now compare the server-side runtime of our MSS protocol against the state-of-the-art PGR scheme by feldman2022. We run both protocols on the Zipf dataset of size n=10,000n=10{,}000 and domain size k=22,000k=22{,}000, across several privacy levels. Table 2 reports the average and standard deviation of the server decoding time (in seconds) over 300 trials. MSS consistently outperforms PGR by large margins, achieving decoding speed-ups between $11\times$ and $448\times$. This performance gap stems from their algorithmic differences: MSS solves a sparse weighted least-squares problem, while PGR relies on algebraic decoding over finite fields. We do not plot GRR here, since its server cost is essentially a single histogram pass $O(n+k)$ and thus serves as a trivial lower bound on runtime; however, as Fig. 1 and Fig. 3 show, GRR is not competitive in our multi-constraint regime due to its much worse utility and attackability. The runtime spike at $\varepsilon=4.5$ for PGR likely arises from parameter rounding and structural constraints in its projective geometry design.
@@
We now evaluate the vulnerability of each LDP protocol to a single-message data reconstruction attack (DRA) (see Section 4.4). Fig. 3 shows the empirical DRA as a function of the privacy budget $\varepsilon$, under the Zipf distribution for domain sizes $k=100$ and k=1,024k=1{,}024. MSS consistently achieves the lowest DRA across all $\varepsilon$ values, confirming its robustness to reconstruction attacks. This is due to its modular randomization strategy, which distributes the probability mass across multiple residue classes, making inference more challenging. In contrast, GRR and SS exhibit higher attackability, especially for small $k$, as their output space is tightly linked to the input domain. PGR behaves comparably to SS/MSS/OUE at moderate $\varepsilon$, but for larger budgets its DRA increases sharply when $k$ is smaller than the projective-domain size $K=(q^{t}-1)/(q-1)$ required by its internal geometry. This truncation mismatch breaks PGR’s combinatorial symmetry and makes certain messages disproportionately informative. A fair comparison using the non-truncated setting $k=K$ is provided in Appendix D.
@@ D Complementary Results
Fig. 4 compares the analytical MSE derived for both SS and MSS against their empirical counterparts, under both Zipf and Spike distributions for $k\in\{1{,}024,22{,}000\}$. The analytical expressions closely match the empirical observations across all regimes of $\varepsilon$, consistently tracking the empirical trends. This validates the tightness and reliability of our closed-form MSE derivation for MSS (Eq. (3)) in practice, regardless of the underlying data distribution or domain size. The results provide strong evidence that our theoretical analysis generalizes well and can be used for practical design decisions without the need for repeated empirical tuning.
@@
Figure 4: Comparison between analytical and empirical MSE for SS and MSS protocols, across a range of privacy budgets $\varepsilon$, for k=1,024k=1{,}024 and k=22,000k=22{,}000 under both Zipf and Spike distributions. Each empirical MSE is averaged over 300 runs, while analytical curves are computed in closed-form expressions.
@@
Fig. 5 presents an ablation study analyzing how the performance of the MSS protocol varies with fixed numbers of moduli $\ell\in\{3,6,9,12,15\}$, compared to the analytically optimized MSS[OPT]. We observe that the relationship between $\ell$, utility (MSE), and communication cost (bit cost) is non-linear: in some regimes, intermediate values such as $\ell=9$ can outperform both smaller ($\ell=3$) and larger ($\ell=12$) settings in terms of accuracy and communication efficiency. This highlights the complex trade-offs induced by modular encoding, where increasing $\ell$ does not guarantee monotonic improvements. Notably, MSS[OPT] consistently selects a configuration that achieves near-optimal utility across all privacy budgets, confirming the effectiveness of our moduli selection strategy. It is important to emphasize that our optimization procedure (Section 4.3 and Appendix A) currently targets minimizing analytical MSE only. This objective implicitly balances communication and robustness in many settings, but it is not guaranteed to yield optimal trade-offs across all criteria. Future work could extend this framework to support multi-objective optimization, for instance, incorporating additional metrics such as bit cost or attackability, enabling more fine-grained control over privacy-utility-efficiency trade-offs.
Figure 5: Ablation study showing the impact of the number of moduli $\ell\in\{3,6,9,12,15\}$ on the utility (left), communication cost (middle), and attackability (right) of the MSS protocol. The dashed black curve represents the performance of our ModularSubsetSelection protocol (i.e., MSS[OPT]), which automatically selects $\ell$ and the moduli via our analytical optimization procedure. Results are averaged over 300 runs for k=1,024k=1{,}024, under the Zipf distribution.
@@
To evaluate the tightness of our analytical DRA derivations (Section 4.4 and Appendix B), we compare theoretical and empirical data reconstruction attack (DRA) for both SS and MSS across privacy budgets. As shown in Fig.6, SS exhibits a near-perfect match between analytical and empirical DRA. For MSS, however, the analytical DRA bound consistently overestimates the true attackability. This gap is theoretically expected: the analytical expression used in the main paper is a conservative upper bound derived using Jensen’s inequality over the random support size of the modular posterior (see Appendix B). The underlying randomness of the filler set in each residue block, and the multiplicity of values per residue, makes the exact computation of MSS’s DRA more intricate, resulting in a looser but guaranteed-safe upper bound. These results show that while the MSS DRA bound is not tight, it still provides a safe analytical proxy and reinforces that MSS offers strong practical protection against reconstruction attacks.
Figure 6: Empirical vs. analytical data reconstruction attack (DRA) under Zipf and Spike distributions, for both small and large domains. While $S$ closely matches its analytical DRA, MSS shows a consistent gap, confirming that the bound is conservative.
@@
While these trends hold generally, it is important to highlight a distinctive behavior of PGR: unlike other mechanisms, PGR’s internal domain size is determined by the projective geometry induced by the privacy budget $K(\varepsilon)=(q^{t}-1)/(q-1)$ with $q\approx e^{\varepsilon}+1$. To isolate the impact of this dependence, we perform an additional ablation (Fig. 9) in which, for each privacy level $\varepsilon$, we set the evaluation domain to the corresponding projective size $K(\varepsilon)$ and run all protocols on this shared domain. We repeat this ablation using $\mathrm{base\_k}=100$ and 1,0241{,}024, to confirm that the phenomenon is consistent across both small and large domains. This geometry-aligned setting removes the truncation mismatch that affects PGR when $k<K(\varepsilon)$ and places all baselines in the domain naturally required by PGR. As shown in Fig. 9, once this alignment is enforced, PGR’s attackability no longer exhibits the sharp rise seen in Figs. 3 and 8, while MSS continues to offer the lowest DRA across all $\varepsilon$.
@@
(a) Spike, $\mathrm{base\_k}=100$.
(b) Spike, $\mathrm{base\_k}=1{,}024$.
Figure 9: Geometry-aligned ablation for PGR. For each privacy level $\varepsilon$, the evaluation domain is set to the projective size $K(\varepsilon)$ induced by PGR, computed using $\mathrm{base\_k}=100$ and 1,0241{,}024. All protocols are evaluated on this shared domain to remove the truncation mismatch that occurs when $k<K(\varepsilon)$. Under this aligned setting, PGR no longer exhibits the high-$\varepsilon$ spike observed in Figs. 3 and 8, while MSS remains the most resistant to reconstruction attacks.